# Yangming Zhang*    Wenqi Jia* Dept. of Computer Science

### v2

@@ 3.4 Overall Workflow of GaussianSpa
We have visualized the evolution of Gaussian opacity distribution along with the PSNR, trained on the Room scene, as shown in Figure 4. It is observed that the opacity distribution of GaussianSpa is significantly distinct from the original 3DGS – there is a clear gap between the “zero" Gaussians and the remaining Gaussians in GaussianSpa, while the PSNR keeps increasing. This means GaussianSpa has successfully imposed a substantial sparsity property and preserved information to the “non-zero" Gaussians. At iteration 25K, we remove all “zero" Gaussians and perform a light tuning that further improves the performance, finally obtaining a compact model with high-quality rendering.
@@ 4.1 Experimental Settings
Implementation Details. We conduct experiments under the same environment specified in the original 3DGS [25] with the PyTorch framework. Our experimental server has two AMD EPYC 9254 CPUs and eight NVIDIA GTX 6000 Ada GPUs. We use the same checkpoint files as Mini-Splatting [20] for training, which are stored before our “optimizing-sparsifying” process to ensure fair comparisons. GaussianSpa starts “optimizing-sparsifying” at iteration 15K and removes “zero" Gaussians at iteration 25K.
@@ 4.2 Quantitative Results
We also plot quality curves with respect to the Gaussian reduction rate on the Room and Kitchen scenes in Figure 7 (more results are in Figure 13 in the Appendix), showing the robustness of GaussianSpa for different numbers of remaining Gaussians. Figure 7 shows that GaussianSpa outperforms the state-of-the-art method, Mini-Splatting [20], by an average of 0.5 dB across multiple Gaussian reduction rates.
Figure 6: Visualization for rendered Gaussian ellipsoids. GaussianSpa adaptively reconstructs high-frequency (e.g., air outlet of the train, complex textures of the carpet) and low-frequency (e.g., large-area sky, simple textures of the carpet) regions using appropriate numbers of Gaussians, respectively, more precisely matching the original textures and shapes, compared to the state-of-the-art Mini-Splatting [20] and the vanilla 3DGS.
@@ 4.3 Visual Quality Results
Figure 7: Quality-Reduction Rate curves on the (left) Kitchen and (right) Room scenes. GaussianSpa outperforms Mini-Splatting [20] across all reduction rates.
As depicted in Figure 5, we present the rendered images on the Drjohnson and Counter scenes [5, 23] for visual quality evaluation, in comparison to LightGaussian [19], Mini-Splattin [20], and the vanilla 3DGS. Together with Figure 1, those outcomes demonstrate the superior rendering quality of our GaussianSpa on diverse scenes over existing approaches, consistent with the quantitative results in the previous subsection. For instance, as shown in Figure 5, GaussianSpa accurately renders the outlet behind the railings for the Playroom scene, while LightGaussian [19] and Mini-Splatting [20] lose this object. Furthermore, GaussianSpa fully recovers the original details of the zoom-in object in the Counter scene, while other methods outcome blurry surfaces.
We comprehensively analyze the proposed GaussianSpa by visualizing the rendered Gaussian ellipsoids and final point clouds in Figure 6 and Figure 8, respectively. Those visualized results demonstrate that GaussianSpa can adaptively control the density to match the details when sparsifying the Gaussians, ensuring high-quality sparse 3D representation. From Figure 6, we can observe that GaussianSpa uses fewer but larger Gaussians to represent the blue sky and uses dense Gaussians to render complex carpet textures, providing precise rendering of details in the zoom-in images. In contrast, Mini-Splatting [20] cannot fully reconstruct the details, leading to blurry rendered textures. Figure 8 further demonstrates our superiority in simplification among existing approaches. GaussianSpa adaptively reduce the Gaussians in the low-frequency areas and preserve more Gaussians in high-frequency areas that generally need more Gaussians to synthesize, resulting in sparse Gaussian point clouds that precisely outline the scene contours.
@@ 5 Discussion
5.1 Convergence Analysis
Figure 8: Visualized point clouds for the Room and Drjohnson scenes. The numbers of remaining Gaussians in millions are displayed. With GaussianSpa, sparse Gaussians concentrate on high-frequency areas after our “optimizing-sparsifying"-integrated training process, ensuring the capability to capture detail-rich textures and shapes. On the contrary, the point clouds produced by other approaches cannot outline the contours.
We analyze convergence behavior of the proposed GaussianSpa and appropriately set hyper-parameters, such as $\delta$, which controls the sparsity strength in Eq. 10. As introduced in Section 3, GaussianSpa performs the “sparsifying" step every fixed number of iterations, referred to here as “interval" here. We plot loss curves for multiple $\delta$’s and intervals in Figure 9. It is seen that with different $\delta$ and interval values, the curves converge well at a similar rate during our “optimizing-sparsifying"-integrated training process. After removing “zero" Gaussians at iteration 25K, the loss curve still exhibits a consistent convergence, confirming the feasibility of GaussianSpa.
Figure 9: Loss curves with multiple (left) $\delta$ and (right) interval settings. The curves show that GaussianSpa exhibits a good convergence behaviour in the “optimizing-sparsifying"-integrated training process.
@@ 5.2 Generality of GaussianSpa
GaussianSpa is a general Gaussian simplification framework. Existing importance measurement criteria can be applied to project the auxiliary variable $\bm{z}$ onto the sparse space in the “sparsifying" step, Eq. 15, obtaining more compact 3DGS models with improved rendering quality. For example, we employ the criterion from LightGaussian [19] and plot the PSNR curves at multiple Gaussian reduction rates in Figure 13 in the Appendix. It is observed that GaussianSpa consistently outperforms LightGaussian [19] with an average of 0.4 dB improvement for the Playroom and Kitchen scenes.
@@ 5.3 Storage Compression
While GaussianSpa focuses on simplifying and optimizing the number of Gaussians to obtain a sparse 3DGS representation, existing compression techniques such as SH optimization and vector quantization can be applied as add-ons to the simplified 3DGS model. As shown in Table 2, we present the final storage cost for the Mip-NeRF 360 dataset by incorporating the SH distillation and vector quantization methods proposed by LightGaussian [19]. This comparison demonstrates that GaussianSpa can achieve the lowest storage consumption with additional compression techniques while maintaining superior rendering quality over existing hybrid compression methods, such as EfficientGS [31] and LightGaussian [19].
@@ 5.4 Relation to Other Optimization Algorithms
The “optimizing-sparsifying" algorithm in our proposed GaussianSpa can be categorized into the class of optimization algorithms that project the target variable onto a constrained space, specifically the $\ell_{p}$ ball (where $p=0$ in our case). This approach is similar to foundational optimization algorithms in compressed sensing [16] and sparse learning [9], such as Ridge regression [24] and Lasso regression [48]. These algorithms extend the least squares regression by incorporating $\ell_{p}$ regularization – Ridge regression adds the $\ell_{2}$ regularization, while the Lasso regression adds the $\ell_{1}$ regularization. The regularization acts as a penalty, combined with quadratic loss, reformulating the problem as a convex optimization task.
@@
However, neither $\ell_{1}$ nor $\ell_{2}$ regularization yields truly sparse solutions; achieving genuine sparsity requires the $\ell_{0}$ regularization. Yet, when $p=0$, the problem, also known as the sparse coding problem [41], becomes nonconvex and is generally recognized as NP-hard [15, 37, 49]. This complexity makes it computationally impractical when the dimension of the target variable is high. In such cases, Lasso (where $p=1$) offers a convex constraint region, serving as the tightest convex surrogate for the sparse coding problem, albeit with some sacrifice in exact sparsity [11]. Another alternative solver suitable for this case is a greedy algorithm, such as the Matching Pursuit algorithm [33, 45]. However, this algorithm cannot guarantee the global optima.
Iterative thresholding (IT) is another direction to address this problem effectively [17, 14, 18]. These algorithms perform a matrix multiplication and its adjoint at each iteration, followed by a scalar shrinkage step, which yields a simple yet highly effective structure. Furthermore, a closed-form solution for the global minimizer of the sparse coding problem can be obtained – even for non-convex cases – under the condition that a unitary matrix is involved [17]. Inspired by this, we introduce a duplicated variable (i.e., $\bm{z}$) and project it onto the $\ell_{0}$ ball, optimizing over two variables alternately, similar to the alternating direction method of multipliers [8]. By leveraging the simplicity of the unitary case, the “sparsifying" step aligns with this scenario and thus admits an explicit solution. Hence, our algorithm simultaneously benefits from the sparsity imposed by the $\ell_{0}$ constraint and achieves fast convergence, guided by the explicit solution.
6 Conclusion
In this paper, we present an optimization-based 3DGS simplification framework, GaussianSpa, for compact and high-quality sparse view synthesis. GaussianSpa formulates 3DGS simplification as a constrained optimization problem with a sparsity constraint on the Gaussian opacities. Then, we propose an “optimizing-sparsifying" solution to efficiently solve the formulated problem during the 3DGS training process. We comprehensively evaluate the proposed GaussianSpa on multiple datasets with quantitative results and qualitative analyses, demonstrating the superiority in rendering quality with fewer Gaussians compared to existing state-of-the-art approaches.
Appendix
7 Additional Results
7.1 Additional Quantitative Results
7.2 Additional Visual Quality Results
Figure 10 and Figure 11 provide additional rendered images comparing GaussianSpa with LightGaussian [19], Mini-Splatting [20], and original 3DGS on various scenes. Those additional results demonstrate our GaussianSpa achieves stronger representational power for background and detail-rich areas such as walls, carpets, and ladders, showcasing superior rendering qualities. We also present additional point cloud views in Figure 12. This further illustrates that GaussianSpa creates a high-quality sparse 3D representation that adaptively uses more Gaussians to represent high-frequency areas.
Scene Method PSNR↑ SSIM↑ LPIPS↓ # G (M)↓
Drjohnson 3DGS 28.77 0.900 0.250 3.260
Mini-Splatting 29.37 0.904 0.261 0.377
GaussianSpa 29.89 0.913 0.243 0.450
GaussianSpa 29.82 0.909 0.254 0.293
Playroom 3DGS 30.07 0.900 0.250 2.290
Mini-Splatting 30.72 0.914 0.248 0.417
GaussianSpa 30.84 0.916 0.254 0.219
GaussianSpa 30.84 0.916 0.254 0.219
Average 3DGS 29.42 0.900 0.250 2.780
Mini-Splatting 30.05 0.909 0.254 0.397
GaussianSpa 30.37 0.914 0.249 0.335
GaussianSpa 30.33 0.912 0.254 0.256
Table 3: Deep Blending per scene results. 3DGS results are reported from [22]. Mini-Splatting [20] results are replicated using official code.
@@ 7.2 Additional Visual Quality Results
Table 5: Tanks&Temples per scene results. 3DGS results are reported from [22]. Mini-Splatting [20] results are replicated using official code.

### v3

@@ 3.4 Overall Workflow of GaussianSpa
We visualize the evolution of Gaussian opacity distribution along with the PSNR, trained on the Room scene, as shown in Figure 5. The opacity distribution of GaussianSpa is significantly distinct from the original 3DGS – there is a clear gap between the “zero" Gaussians and the remaining Gaussians in GaussianSpa, while the PSNR keeps increasing. This means GaussianSpa has successfully imposed a substantial sparsity property and preserved information to the “non-zero" Gaussians. At iteration 25K, we remove all “zero" Gaussians and perform a light tuning that further improves the performance, finally obtaining a compact model with high-quality rendering.
@@ 4 Experiments
Figure 6: Visualization for rendered Gaussian ellipsoids. (Add’l Figure 12 in the Supplementary Material). GaussianSpa renders high-frequency (e.g., air outlet of the train) and low-frequency (e.g., large-area sky) areas adaptively using appropriate numbers of Gaussians, respectively, more precisely matching the original textures, compared to the state-of-the-art Mini-Splatting [17] and the vanilla 3DGS.
@@ 4.1 Experimental Settings
Implementation Details. We conduct experiments under the same environment specified in the original 3DGS [22] with the PyTorch framework. Our experimental server has two AMD EPYC 9254 CPUs and eight NVIDIA GTX 6000 Ada GPUs. We use the same starting checkpoint files as Mini-Splatting [17] before our “optimizing-sparsifying” process to ensure fair comparisons. GaussianSpa starts “optimizing-sparsifying” at iteration 15K and removes “zero" Gaussians at iteration 25K.
Figure 7: Quality-Reduction Rate curves on the (left) Kitchen and (right) Room scenes. GaussianSpa outperforms Mini-Splatting [17] at multiple reduction rates.
@@ 4.2 Quantitative Results
We plot quality curves for the Gaussian reduction rate on the Room and Kitchen in Figure 7 (more results are in Figure 14 in the Supplementary Material), showing the robustness of GaussianSpa for different numbers of remaining Gaussians. GaussianSpa outperforms the state-of-the-art method, Mini-Splatting [17], by an average of 0.5 dB PSNR improvement across multiple Gaussian reduction rates.
Figure 8: Visualized point clouds for the Room and Drjohnson scenes. With GaussianSpa, sparse Gaussians concentrate on high-frequency areas after our “optimizing-sparsifying"-integrated training process, ensuring the capability to capture detail-rich textures and shapes. On the contrary, the point clouds produced by other approaches cannot outline the contours.
@@ 4.3 Visual Quality Results
Figure 4 shows the rendered images on the Drjohnson and Counter scenes [3, 20] for visual quality evaluation, in comparison to LightGaussian [16], Mini-Splatting [17], and the vanilla 3DGS. Together with Figure 1, these outcomes demonstrate GaussianSpa’s superior rendering quality on diverse scenes over existing approaches, consistent with the quantitative results in the previous subsection. For instance, as shown in Figure 4, GaussianSpa accurately renders the outlet behind the railings for the Playroom scene, while LightGaussian [16] and Mini-Splatting [17] lose this object. Furthermore, GaussianSpa fully recovers the original details of the zoom-in object in the Counter scene, while other methods outcome blurry surfaces.
We analyze the proposed GaussianSpa by visualizing the rendered Gaussian ellipsoids and final point clouds in Figure 6 and Figure 8, respectively. Those visualized results demonstrate that GaussianSpa can adaptively control the density to match the details when sparsifying the Gaussians, ensuring high-quality sparse 3D representation. From Figure 6 and Figure 12, we can observe that GaussianSpa uses fewer but larger Gaussians to represent the blue sky and uses dense Gaussians to render complex carpet textures, providing precise rendering of details in the zoom-in images. In contrast, Mini-Splatting [17] cannot fully present the details, leading to blurry rendered textures. Figure 8 further demonstrates the superiority of GaussianSpa in simplification over existing methods. GaussianSpa adaptively reduce the Gaussians in the low-frequency areas and preserve more Gaussians in high-frequency areas that generally need denser Gaussians to synthesize, resulting in sparse Gaussian point clouds that precisely outline the scene contours.
@@ 5.1 Generality of GaussianSpa
The proposed GaussianSpa is a general Gaussian simplification framework. Existing importance criteria can be applied to project the auxiliary variable $\bm{z}$ onto the sparse space in the “sparsifying" step, Eq. 15, obtaining more compact 3DGS models with improved rendering performance. For example, we employ the criterion from LightGaussian [16] and plot the PSNR curves at multiple Gaussian reduction rates in Figure 14 in the Supplementary Material. It is observed that GaussianSpa consistently outperforms LightGaussian [16] with an average of 0.4 dB improvement for the Playroom and Kitchen scenes.
@@ 5.2 Storage Compression
While GaussianSpa focuses on simplifying and optimizing the number of Gaussians for a sparse 3DGS representation, existing compression techniques such as SH optimization and vector quantization can be applied as add-ons to the simplified 3DGS model. Table 2 (in Supplementary Material) shows the final storage cost for the Mip-NeRF 360 dataset by incorporating the SH distillation and vector quantization methods proposed by LightGaussian [16]. This comparison demonstrates that GaussianSpa can achieve the lowest storage consumption while maintaining superior rendering quality over existing hybrid compression methods, such as EfficientGS [28] and LightGaussian [16].
@@ 5.3 Relation to Other Optimization Algorithms
The “optimizing-sparsifying" solution in our proposed GaussianSpa falls under optimization methods that project a target variable onto a constrained space, specifically the $\ell_{p}$ ball (where $p=0$ in our case). This approach is similar to foundational optimization algorithms in compressed sensing [13] and sparse learning [6], such as Ridge regression [21] with the $\ell_{2}$ regularization and Lasso regression [45] with the $\ell_{1}$ regularization. The regularization acts as a penalty, combined with quadratic loss, reformulating the problem as a convex optimization task.
However, genuine sparsity requires $\ell_{0}$ regularization, which leads to a nonconvex, NP-hard problem [38, 12, 34, 46]. Lasso (where $p=1$) provides the tightest convex approximation but sacrifices exact sparsity [8]. Greedy algorithms like Matching Pursuit [30, 42] offer another approach, though they lack guaranteed global optima.
Iterative thresholding (IT) [14, 11, 15] offers a different solution, using matrix multiplications and scalar shrinkage steps to yield a simple yet effective structure. Furthermore, a closed-form solution for the global minimizer can be obtained in cases with a unitary matrix [14], even for non-convex scenarios. Inspired by this, we introduce a duplicated variable (i.e., $\bm{z}$) and project it onto the $\ell_{0}$ ball, optimizing over two variables alternately, similar to the alternating direction method of multipliers [5]. Leveraging the unitary case’s simplicity, the “sparsifying" step aligns with this scenario and thus admits an explicit solution. Thus, our algorithm benefits from $\ell_{0}$-induced sparsity while achieving fast convergence guided by the explicit solution.
6 Conclusion
In this paper, we present an optimization-based simplification framework, GaussianSpa, for compact and high-quality 3DGS. GaussianSpa formulates the simplification objective as a constrained optimization problem with a sparsity constraint on the Gaussian opacities. Then, we propose an “optimizing-sparsifying" solution to efficiently solve the formulated problem during the 3DGS training process. Our comprehensive evaluations on multiple datasets with quantitative results and qualitative analyses demonstrate the superiority of GaussianSpa in rendering quality with fewer Gaussians compared to the state of the arts.
Supplementary Material
7 Additional Results
7.1 Additional Quantitative Results
7.2 Additional Visual Quality Results
Figure 10 and Figure 11 provide additional rendered images comparing GaussianSpa with LightGaussian [16], Mini-Splatting [17], and original 3DGS on various scenes. Those additional results demonstrate our GaussianSpa achieves stronger representational power for background and detail-rich areas such as walls, carpets, and ladders, showcasing superior rendering qualities. Furthermore, Figure 12 and Figure 13 offer additional rendered Gaussian ellipsoids and point cloud views, respectively. These results further illustrate that GaussianSpa creates a high-quality sparse 3D representation that adaptively uses more Gaussians to represent high-frequency areas.
8 Additional Discussion
@@
Convergence Analysis. We analyze the convergence behavior of GaussianSpa by examining the effects of hyper-parameters such as $\delta$, which controls the sparsity strength in Eq. 10, and the interval at which the “sparsifying" step is performed, as described in Section 3. Figure 9 shows loss curves for various $\delta$ and interval values, indicating similar convergence rates during our “optimizing-sparsifying"-integrated training process. After removing “zero" Gaussians at iteration 25K, the loss curve continues to converge consistently, confirming GaussianSpa’s feasibility.
Figure 9: Loss curves with multiple (left) $\delta$ and (right) interval settings. The curves show that GaussianSpa exhibits a good convergence behavior in the “optimizing-sparsifying"-integrated training process.
@@
Figure 12: Additional visualized Gaussian ellipsoids. The numbers of remaining Gaussians in millions are displayed.
@@
Table 4: Tanks&Temples per scene results.
Scene Method PSNR↑ SSIM↑ LPIPS↓ # G (M)↓
Drjohnson 3DGS 28.77 0.900 0.250 3.260
Mini-Splatting 29.37 0.904 0.261 0.377
GaussianSpa 29.89 0.913 0.243 0.450
GaussianSpa 29.82 0.909 0.254 0.293
Playroom 3DGS 30.07 0.900 0.250 2.290
Mini-Splatting 30.72 0.914 0.248 0.417
GaussianSpa 30.84 0.916 0.254 0.219
GaussianSpa 30.84 0.916 0.254 0.219
Average 3DGS 29.42 0.900 0.250 2.780
Mini-Splatting 30.05 0.909 0.254 0.397
GaussianSpa 30.37 0.914 0.249 0.335
GaussianSpa 30.33 0.912 0.254 0.256
Table 5: Deep Blending per scene results.